A meat market manager mixes hamburger having 30% fat content with hamburger that has 10% fat content in order to obtain 400 pounds of hamburger with 25% fat content. How much hamburger of each type should the manager use?
Answer by nyc_function(2741)

30x + 10y = 25( x + y)
6x + 2y = 5x + 5y
x = 3y, or 3 lb of x (30% fat) for each pound of y (10% fat)
300 lb 30% + 100 lb 10%
I wouldn't buy a hamburger here!